16. Give the number of protons, electrons, and neutrons in each of the following: 10 a. c. 6329 Cu d. 7734 Se 17. Write the symbol for each of the following ions: a. the ion with 3+ charge, 28 electrons, and a mass number of 71 b. the ion with 36 electrons, 35 protons, and 45 neutrons c. the ion with 86 electrons, 142 neutrons, and a 4+ charge d. the ion with a 2+ charge, atomic number 38, and mass number 87 18. Determine the empirical formula. Acetic acid, C2H4O2 b. Citric acid, C&HsO c. Hydrazine, N2H4 d. Nicotine, CioHuN2 e. Butane, C4Hio a. 19. What is the difference between atomic mass, mass number, and average mass?
Answer #1

16)

a)

number of ptoton is number at bottom which is 5

SO,

number of proton = 5

Since number of proton = number of electron,

number of electron = 5

number of neutron = mass number - number of proton = 10 - 5 = 5

b)

number of proton = 80

number of electron = 80

number of neutron = 199 - 80 = 119

c)

number of proton = 29

number of electron = 29

number of neutron = 63 - 29 = 34

d)

number of proton = 34

number of electron = 34

number of neutron = 77 - 34 = 43

I am allowed to answer only 1 question at a time